# Grounded in Literature

### v1

@@ A.2.2 Prompts
Dimension Extraction Prompt You are a helpful assistant that reads scientific research ideas and extracts a structured summary of their contributions. Your task is to identify a small number of high-level contribution dimensions, and for each dimension, extract one or more specific contribution statements that are faithful to the research idea. Contribution dimensions should represent general categories of scientific contribution that are meaningful and comparable across research ideas, regardless of the field. These might include: - methodology (e.g., proposing a new method, model, or procedure) - application (e.g., applying existing methods to a new problem or domain) - theoretical contribution (e.g., proving a new result, deriving a new model) - data (e.g., constructing a new dataset, conducting original measurements or surveys) - evaluation (e.g., designing an experimental protocol, benchmarking a technique) - tool or system design (e.g., building software, devices, or infrastructure to support research) - conceptual framework (e.g., introducing a new taxonomy or way of thinking about a problem) Do not limit your output to the examples above but rather generate suitable dimensions for the research idea given to you. Only include dimensions that are actually reflected in the research idea — do not add generic or speculative categories. For each dimension, write one or more contribution statements that clearly explain what the research is proposing. These statements should be precise, self-contained, and informative — make sure they include enough context as they will be used as the basis to generate search queries later on. Each dimension may include multiple contribution statements, but do not repeat the same idea across dimensions. Avoid redundancy, and keep the summary compact and informative. Please output a parseable JSON block as follows:
‘‘‘json
  "<dimension_name_1>": [
    "<contribution_statement_1>",
    "<contribution_statement_2>"
  "<dimension_name_2>": [
    "<contribution_statement_3>"
@@
Relevance Assessment Prompt You are an expert at evaluating whether a paper should be considered relevant for assessing the scientific contribution of a research idea. You will be given a research idea and a paper abstract retrieved from Semantic Scholar. Your task is to thoroughly understand the scientific contributions of each of the research idea and the paper, and to output a score from 0 to 5 indicating how similar the contributions of the paper are to those of the research idea. Scoring Rubric: Score 5 — Highly Similar Contributions: The paper addresses the exact same research question or hypothesis as the idea Uses identical or very similar methodological approaches Targets the same specific population, system, or domain Would directly compete with or overlap significantly with the proposed research The paper’s findings would substantially impact the novelty of the proposed work Score 4 — Very Similar Contributions: The paper addresses a closely related research question with significant overlap Uses similar methodological approaches with minor variations Targets a very similar population, system, or domain Shares most key variables, measurements, or outcomes of interest The paper’s contributions would moderately impact the proposed research’s novelty Score 3 — Moderately Similar Contributions: The paper addresses a related research question within the same broad area Uses some similar methods or approaches but with notable differences Targets a related but distinct population, system, or domain Shares some key concepts, variables, or theoretical frameworks The paper provides useful context but doesn’t directly threaten novelty Score 2 — Somewhat Similar Contributions: The paper is in the same general field or discipline Uses different methods but addresses conceptually related problems Limited overlap in specific research focus or target populations Shares broad theoretical background but differs in specific contributions The paper is peripherally relevant for background or context Score 1 — Minimally Similar Contributions: The paper is tangentially related to the research area Very limited overlap in methods, populations, or specific research questions May share some terminology or broad field classification Provides minimal insight relevant to the proposed research Connection is primarily at the disciplinary level Score 0 — No Similar Contributions: The paper addresses completely different research questions No meaningful overlap in methods, populations, or domains Different field or discipline entirely No relevant insights for the proposed research No discernible connection between the contributions Respond with a JSON object containing two fields rationale and score. Please output a parseable JSON block as follows:
‘‘‘json
  "rationale": "<your detailed reasoning for the score>",
  "score": <an integer from 0 to 5 based on the rubric>
@@
Pairwise Comparison Prompt You are an expert in assessing the novelty of scientific research ideas relative to existing work. Your job is to compare a full research idea to a paper abstract based on novelty and originality. You will be given a research idea, a paper abstract, and a comma-separated list of contribution dimensions. Produce a structured output consisting of: 1. overall_comparison: a broad yet precise summary of the novelty of the research idea versus the paper, i.e., whether the idea proposes ideas/angles/uses that appear original relative to what the abstract claims; identify overlap vs. originality explicitly. 2. dimension_comparisons: for each provided dimension, a novelty comparison that states whether the idea is doing something not present in the paper for that dimension (or vice versa), and a numeric score: 1 = The research idea is more novel under this dimension (adds ideas/angles/uses not present in the paper abstract) 0 = Neither appears more novel or the paper does not address this dimension (tie/insufficient evidence) -1 = The paper appears more novel or the idea largely replicates what the paper already presents under this dimension Return your output as a parseable JSON block with exactly this structure:
‘‘‘json
  "overall_comparison": "<novelty-focused summary>",
  "dimension_comparisons": {
    "<dimension_1>": {
      "comparison": "<comparison for this dimension>",
      "score": <1 | 0 | -1>
    "<dimension_2>": {
      "comparison": "<comparison>",
      "score": <1 | 0 | -1>
@@
Appendix B ScholarEval interface
@@
Figure 5: ScholarEval user interface.
Appendix C ScholarIdeas curation details
C.1 Paper List
We provide the full list of 117 papers used to construct ScholarIdeas, at the end of the manuscript in LABEL:tab:ai-paper-list, LABEL:tab:neuro-paper-list, LABEL:tab:biochem-paper-list, and LABEL:tab:eco-paper-list. Each entry in the table links to the paper on OpenReview or eLife.
C.2 LLM-based extraction
C.2.1 Details
We use Claude 4 Sonnet to extract the research idea and review rubrics. The extraction of the review rubrics is done in two stages: (1) we instruct the LLM to extract the verbatim excerpts from the review that give assessments about the research idea. (2) we instruct the LLM to remove redundancies and format the extracted text into standalone statements classified based on type, axis, and severity. The full prompts used in this process are in §C.2.2.
C.2.2 Prompts
@@
Verbatim Review Extraction You are an expert assistant skilled at extracting key components of scientific paper reviews containing evaluations of the underlying research idea and discarding all other criticism of other aspects of the scientific paper, including results and presentation. Namely, you will be given a text containing all the reviews that the scientific paper received as well as a research idea detailing the main research idea contained in the scientific paper, and your job is to extract sentences from the review that address components mentioned in the research idea. Any sentences that contain criticism of aspects of the paper not mentioned in the research idea should not be extracted. This is an important guideline to follow. Reviews contain the following components: Summary: A brief overview of the paper’s contributions, main ideas, and objectives. Strengths: Positive aspects highlighted by the reviewer, such as innovative approaches, solid methodology, or valuable contributions. Weaknesses: Negative aspects or concerns raised by the reviewer, such as lack of clarity, methodological flaws, or unanswered questions. Questions: Queries or points of clarification raised by the reviewer, often regarding experimental design, assumptions, or aspects of the research idea. The research idea corresponding to the research paper includes: Problem: Background information on the problem the paper addresses, the motivation for pursuing it, and any hypotheses formed by the authors. Method: The methodology proposed by the authors to address the problem. Experiment Design: Details on the experiments, data collection, and evaluation techniques that the authors plan to employ. Review:
{review_text}
Research Idea:
{research_idea_text}
Output format: This is an extraction task, not a generation task. Your job is exactly to extract sentences from the reviews which address the ideation behind the research (i.e., the research at the idea stage, imagine that no results have been obtained or paper written yet). These should only include criticism related to the problem that the scientists are trying to solve, the methodology they are employing, or the experiments they plan to conduct. Any other content related to typographical errors, presentation or formatting issues, results, comments on figures and tables, artifacts, or conclusions derived from executing the experiments should absolutely not be included. Only points related to the underlying research idea as specified in the research idea should be kept. Any point not related to an aspect explicitly mentioned in the research idea should not be included. Return the extracted sentences (verbatim! no changes made) as follows:
Extracted Review: <verbatim extracted sentences>
Nothing else besides the extracted review should be contained in the response.
@@
Review Formatting Into Rubrics You are an expert assistant skilled at extracting clear and non-redundant evaluation statements from a review summary. In this task you will be given a summary that was created by merely concatenating extracted sentences from a longer review. You will be given this extractive review summary along with the research idea that it reviews and you have to do the following tasks while taking into account the constraints. Tasks: Removing redundancies: The review summary was created by concatenating the extracted sentences from multiple reviews written by different reviewers about the same paper. Thus, some of the sentences can be redundant. You should remove all redundancies. If a point is mentioned more than once, drop the other mentions and only keep one. Express each evaluation (weakness or strength mentioned by the reviewer) as a clear statement. Hence your final output should be a list of such statements, each one addressing a specific point made by the reviewer(s). Adding context: The review summary might refer to some aspects of the research idea without sufficient context. You should aim to make each statement in the list clearer by adding context from the research idea that would enhance the understanding of the review. Remove mentions of the final paper: These statements are extracted from the reviews of a paper submitted to a conference or journal. Hence, some statements might have mentions of the final paper (e.g., “in line so and so …” or “in Figure x …”). When such statements are encountered, you should slightly rephrase them — without alterations to the meaning — to ensure that they do not refer to the paper but rather to the corresponding component in the research idea. Ensuring this consistency between the statements and the research idea content is crucial. For each statement (item in the list), you should also annotate it with whether it is a "strength" or a "weakness", whether it is related to the paper’s "soundness" (correctness of the methods, etc.) or "contribution" (novelty and contribution to the field), and finally annotate whether it represents a "major" or "minor" point. Major points are those that reviewers highly stress and that affect the overall evaluation of the paper, while minor points are those that are mentioned but not stressed as much. Constraints: Make sure that the changes you make are minimal and only aim to remove redundancy, express the statements as a list, add context, and ensure consistency with the research idea content. Do not completely paraphrase the review summary. Try to keep the original wording and sentences as much as possible. Your changes should not lead to a significant overhaul of the review summary. Do not change the tone of the review. It should still be written as if produced by an actual reviewer. Extractive review summary:
{intermediate_text}
Research Idea:
{research_idea_text}
Please output a parseable JSONL (JSON Lines) block where each item in the JSONL is a JSON object on a single line as follows:
‘‘‘jsonl
  "statement": "<one clear, self-contained statement>",
  "type": "strength",
  "axis": "soundness",
  "severity": "major"
  "statement": "<another clear, self-contained statement>",
  "type": "weakness",
  "axis": "contribution",
  "severity": "minor"
  "statement": "<another clear, self-contained statement>",
  "type": "strength",
  "axis": "soundness",
  "severity": "minor"
C.3 Expert Validation
C.3.1 Details
@@
We invited six experts of Ph.D. students, postdoctoral researchers, and professors (1 from computer science, 2 from biochemistry, 2 from neuroscience, and 1 from ecology). We first conducted a training session to explain the validation task and we have provided the annotators with detailed instructions, shown in §C.3.2. We made an effort to match every annotator to research ideas in their specific area of specialty, and annotators were instructed to skip a research idea if they felt that it fell out of their area of expertise.
C.3.2 Validation Instructions
@@ C.3.3 Validation Results
Appendix D Evaluation metrics details
D.1 Coverage Metric
@@
Coverage rubric given to Prometheus
rubric_data = {
  "criteria":
  "You are given a single essential statement from
  a research idea review targeting the technical
  soundness and scientific contribution of a research
  idea. This statement was extracted from actual human
  reviews about the research idea and will be used
  as your ground-truth reference. Additionally, you
  will be given a research idea review that aims to
  give a detailed evaluation of the same research idea.
  The review given to you will typically be long and
  composed of multiple sections and references to related
  papers. Your job is to verify that the given review
  has a mention of the reference statement in its content.
  These references can be paraphrased or articulated
  differently, but they should express the same meaning
  as the reference statement. Does the given review
  express the reference statement?",
  "score1_description":
  "The given review does not mention the reference
  statement at all",
  "score2_description":
  "The given review mentions some related concepts
  but does not express the core meaning of the
  reference statement",
  "score3_description":
  "The given review partially expresses the reference
  statement but misses key aspects or nuances",
  "score4_description":
  "The given review expresses most aspects of the
  reference statement with minor omissions or slight
  differences in emphasis",
  "score5_description":
  "The given review clearly expresses the reference
  statement, capturing its core meaning and implications"
D.2 Reference Invalidity
We compute reference invalidity as the fraction of non-resolving citations by issuing an HTTP HEAD request for every paper link referenced in the output of ScholarEval and baselines, and inspecting the returned status code. In our initial implementation, we noticed that this approach undercounts failures: many publishers return 403 (bot blocking) or even 200 (soft-404 pages etc.) for URLs that do not actually resolve to the referenced papers upon manual inspection. We thus resort to reporting a lower bound for reference invalidity; we label a link as invalid only when it returns 404, 410, or any 5xx status. However the true of reference invalidity for baseline systems is higher. We provide failure cases observed in our manual audit in §F.2. Reference invalidity is entirely eliminated in ScholarEval.
D.3 LLM Metrics

### v2

@@ B.2.2 Prompts
Dimension Extraction Prompt You are a helpful assistant that reads scientific research ideas and extracts a structured summary of their contributions. Your task is to identify a small number of high-level contribution dimensions, and for each dimension, extract one or more specific contribution statements that are faithful to the research idea. Contribution dimensions should represent general categories of scientific contribution that are meaningful and comparable across research ideas, regardless of the field. These might include: - methodology (e.g., proposing a new method, model, or procedure) - application (e.g., applying existing methods to a new problem or domain) - theoretical contribution (e.g., proving a new result, deriving a new model) - data (e.g., constructing a new dataset, conducting original measurements or surveys) - evaluation (e.g., designing an experimental protocol, benchmarking a technique) - tool or system design (e.g., building software, devices, or infrastructure to support research) - conceptual framework (e.g., introducing a new taxonomy or way of thinking about a problem) Do not limit your output to the examples above but rather generate suitable dimensions for the research idea given to you. Only include dimensions that are actually reflected in the research idea — do not add generic or speculative categories. For each dimension, write one or more contribution statements that clearly explain what the research is proposing. These statements should be precise, self-contained, and informative — make sure they include enough context as they will be used as the basis to generate search queries later on. Each dimension may include multiple contribution statements, but do not repeat the same idea across dimensions. Avoid redundancy, and keep the summary compact and informative. Please output a parseable JSON block as follows: ‘‘‘json { "<dimension_name_1>": [ "<contribution_statement_1>", "<contribution_statement_2>" ], "<dimension_name_2>": [ "<contribution_statement_3>" ] } ‘‘‘
@@
Relevance Assessment Prompt You are an expert at evaluating whether a paper should be considered relevant for assessing the scientific contribution of a research idea. You will be given a research idea and a paper abstract retrieved from Semantic Scholar. Your task is to thoroughly understand the scientific contributions of each of the research idea and the paper, and to output a score from 0 to 5 indicating how similar the contributions of the paper are to those of the research idea. Scoring Rubric: Score 5 — Highly Similar Contributions: The paper addresses the exact same research question or hypothesis as the idea Uses identical or very similar methodological approaches Targets the same specific population, system, or domain Would directly compete with or overlap significantly with the proposed research The paper’s findings would substantially impact the novelty of the proposed work Score 4 — Very Similar Contributions: The paper addresses a closely related research question with significant overlap Uses similar methodological approaches with minor variations Targets a very similar population, system, or domain Shares most key variables, measurements, or outcomes of interest The paper’s contributions would moderately impact the proposed research’s novelty Score 3 — Moderately Similar Contributions: The paper addresses a related research question within the same broad area Uses some similar methods or approaches but with notable differences Targets a related but distinct population, system, or domain Shares some key concepts, variables, or theoretical frameworks The paper provides useful context but doesn’t directly threaten novelty Score 2 — Somewhat Similar Contributions: The paper is in the same general field or discipline Uses different methods but addresses conceptually related problems Limited overlap in specific research focus or target populations Shares broad theoretical background but differs in specific contributions The paper is peripherally relevant for background or context Score 1 — Minimally Similar Contributions: The paper is tangentially related to the research area Very limited overlap in methods, populations, or specific research questions May share some terminology or broad field classification Provides minimal insight relevant to the proposed research Connection is primarily at the disciplinary level Score 0 — No Similar Contributions: The paper addresses completely different research questions No meaningful overlap in methods, populations, or domains Different field or discipline entirely No relevant insights for the proposed research No discernible connection between the contributions Respond with a JSON object containing two fields rationale and score. Please output a parseable JSON block as follows: ‘‘‘json { "rationale": "<your detailed reasoning for the score>", "score": <an integer from 0 to 5 based on the rubric> } ‘‘‘
@@
Pairwise Comparison Prompt You are an expert in assessing the novelty of scientific research ideas relative to existing work. Your job is to compare a full research idea to a paper abstract based on novelty and originality. You will be given a research idea, a paper abstract, and a comma-separated list of contribution dimensions. Produce a structured output consisting of: 1. overall_comparison: a broad yet precise summary of the novelty of the research idea versus the paper, i.e., whether the idea proposes ideas/angles/uses that appear original relative to what the abstract claims; identify overlap vs. originality explicitly. 2. dimension_comparisons: for each provided dimension, a novelty comparison that states whether the idea is doing something not present in the paper for that dimension (or vice versa), and a numeric score: 1 = The research idea is more novel under this dimension (adds ideas/angles/uses not present in the paper abstract) 0 = Neither appears more novel or the paper does not address this dimension (tie/insufficient evidence) -1 = The paper appears more novel or the idea largely replicates what the paper already presents under this dimension Return your output as a parseable JSON block with exactly this structure: ‘‘‘json { "overall_comparison": "<novelty-focused summary>", "dimension_comparisons": { "<dimension_1>": { "comparison": "<comparison for this dimension>", "score": <1 | 0 | -1> }, "<dimension_2>": { "comparison": "<comparison>", "score": <1 | 0 | -1> } } } ‘‘‘
@@
Appendix C ScholarEval interface
Figure 4: ScholarEval user interface.
@@ Appendix C ScholarEval interface
Appendix D ScholarIdeas curation details
D.1 Paper List
We provide the full list of the 117 papers that were used to construct ScholarIdeas at the end of the manuscript. Each entry in the lists links to the paper on OpenReview or eLife.
D.2 LLM-based extraction
D.2.1 Details
For each paper in our dataset, we use Claude 4 Sonnet to extract the research idea and review rubrics. The extraction of the review rubrics is done in two stages: first, we instruct the LLM to extract the verbatim excerpts from the review that give assessments about the research idea. In the second step, we instruct the LLM to remove redundancies and format the extracted text into standalone statements classified based on type, axis, and severity. The full prompts used in this process are provided in section D.2.2.
D.2.2 Prompts
@@
Verbatim Review Extraction You are an expert assistant skilled at extracting key components of scientific paper reviews containing evaluations of the underlying research idea and discarding all other criticism of other aspects of the scientific paper, including results and presentation. Namely, you will be given a text containing all the reviews that the scientific paper received as well as a research idea detailing the main research idea contained in the scientific paper, and your job is to extract sentences from the review that address components mentioned in the research idea. Any sentences that contain criticism of aspects of the paper not mentioned in the research idea should not be extracted. This is an important guideline to follow. Reviews contain the following components: Summary: A brief overview of the paper’s contributions, main ideas, and objectives. Strengths: Positive aspects highlighted by the reviewer, such as innovative approaches, solid methodology, or valuable contributions. Weaknesses: Negative aspects or concerns raised by the reviewer, such as lack of clarity, methodological flaws, or unanswered questions. Questions: Queries or points of clarification raised by the reviewer, often regarding experimental design, assumptions, or aspects of the research idea. The research idea corresponding to the research paper includes: Problem: Background information on the problem the paper addresses, the motivation for pursuing it, and any hypotheses formed by the authors. Method: The methodology proposed by the authors to address the problem. Experiment Design: Details on the experiments, data collection, and evaluation techniques that the authors plan to employ. Review: {review_text} Research Idea: {research_idea_text} Output format: This is an extraction task, not a generation task. Your job is exactly to extract sentences from the reviews which address the ideation behind the research (i.e., the research at the idea stage, imagine that no results have been obtained or paper written yet). These should only include criticism related to the problem that the scientists are trying to solve, the methodology they are employing, or the experiments they plan to conduct. Any other content related to typographical errors, presentation or formatting issues, results, comments on figures and tables, artifacts, or conclusions derived from executing the experiments should absolutely not be included. Only points related to the underlying research idea as specified in the research idea should be kept. Any point not related to an aspect explicitly mentioned in the research idea should not be included. Return the extracted sentences (verbatim! no changes made) as follows: Extracted Review: <verbatim extracted sentences> Nothing else besides the extracted review should be contained in the response.
@@
Review Formatting Into Rubrics You are an expert assistant skilled at extracting clear and non-redundant evaluation statements from a review summary. In this task you will be given a summary that was created by merely concatenating extracted sentences from a longer review. You will be given this extractive review summary along with the research idea that it reviews and you have to do the following tasks while taking into account the constraints. Tasks: Removing redundancies: The review summary was created by concatenating the extracted sentences from multiple reviews written by different reviewers about the same paper. Thus, some of the sentences can be redundant. You should remove all redundancies. If a point is mentioned more than once, drop the other mentions and only keep one. Express each evaluation (weakness or strength mentioned by the reviewer) as a clear statement. Hence your final output should be a list of such statements, each one addressing a specific point made by the reviewer(s). Adding context: The review summary might refer to some aspects of the research idea without sufficient context. You should aim to make each statement in the list clearer by adding context from the research idea that would enhance the understanding of the review. Remove mentions of the final paper: These statements are extracted from the reviews of a paper submitted to a conference or journal. Hence, some statements might have mentions of the final paper (e.g., “in line so and so …” or “in Figure x …”). When such statements are encountered, you should slightly rephrase them — without alterations to the meaning — to ensure that they do not refer to the paper but rather to the corresponding component in the research idea. Ensuring this consistency between the statements and the research idea content is crucial. For each statement (item in the list), you should also annotate it with whether it is a "strength" or a "weakness", whether it is related to the paper’s "soundness" (correctness of the methods, etc.) or "contribution" (novelty and contribution to the field), and finally annotate whether it represents a "major" or "minor" point. Major points are those that reviewers highly stress and that affect the overall evaluation of the paper, while minor points are those that are mentioned but not stressed as much. Constraints: Make sure that the changes you make are minimal and only aim to remove redundancy, express the statements as a list, add context, and ensure consistency with the research idea content. Do not completely paraphrase the review summary. Try to keep the original wording and sentences as much as possible. Your changes should not lead to a significant overhaul of the review summary. Do not change the tone of the review. It should still be written as if produced by an actual reviewer. Extractive review summary: {intermediate_text} Research Idea: {research_idea_text} Please output a parseable JSONL (JSON Lines) block where each item in the JSONL is a JSON object on a single line as follows: ‘‘‘jsonl { "statement": "<one clear, self-contained statement>", "type": "strength", "axis": "soundness", "severity": "major" } { "statement": "<another clear, self-contained statement>", "type": "weakness", "axis": "contribution", "severity": "minor" } { "statement": "<another clear, self-contained statement>", "type": "strength", "axis": "soundness", "severity": "minor" }
D.3 Expert Validation
D.3.1 Details
@@
We invited six experts of Ph.D. students, postdoctoral researchers, and professors (1 from computer science, 2 from biochemistry, 2 from neuroscience, and 1 from ecology). We first conducted a training session to explain the validation task and we have provided the annotators with detailed instructions, shown in section D.3.2. We made an effort to match every annotator to research ideas in their specific area of specialty, and annotators were instructed to skip a research idea if they felt that it fell out of their area of expertise.
D.3.2 Validation Instructions
@@ D.3.3 Validation Results
Appendix E Evaluation metrics details
E.1 Coverage Metric
@@
Coverage rubric given to Prometheus rubric_data = { "criteria": "You are given a single essential statement from a research idea review targeting the technical soundness and scientific contribution of a research idea. This statement was extracted from actual human reviews about the research idea and will be used as your ground-truth reference. Additionally, you will be given a research idea review that aims to give a detailed evaluation of the same research idea. The review given to you will typically be long and composed of multiple sections and references to related papers. Your job is to verify that the given review has a mention of the reference statement in its content. These references can be paraphrased or articulated differently, but they should express the same meaning as the reference statement. Does the given review express the reference statement?", "score1_description": "The given review does not mention the reference statement at all", "score2_description": "The given review mentions some related concepts but does not express the core meaning of the reference statement", "score3_description": "The given review partially expresses the reference statement but misses key aspects or nuances", "score4_description": "The given review expresses most aspects of the reference statement with minor omissions or slight differences in emphasis", "score5_description": "The given review clearly expresses the reference statement, capturing its core meaning and implications" }
E.2 Reference Invalidity
We compute reference invalidity as the fraction of non-resolving citations by issuing an HTTP HEAD request for every paper link referenced in the output of ScholarEval and baselines, and inspecting the returned status code. In our initial implementation, we noticed that this approach undercounts failures: many publishers return 403 (bot blocking) or even 200 (soft-404 pages etc.) for URLs that do not actually resolve to the referenced papers upon manual inspection. We thus resort to reporting a lower bound for reference invalidity; we label a link as invalid only when it returns 404, 410, or any 5xx status. However the true of reference invalidity for baseline systems is higher. We provide failure cases observed in our manual audit in §G.2. Reference invalidity is entirely eliminated in ScholarEval.
E.3 LLM Metrics